OpenStudy (anonymous):

help i will fan and medal answer already solved need to figure out out how she got the numbers

OpenStudy (anonymous):

I sort of said it wrong it's not slope I always just got to slope be you would use pathegorean theorem for the two sides so 3^2+4^2=c^2 so the length would be root 13 so the perimeter would be root 13 + root 13 +6+6 12+2sqrt13

OpenStudy (anonymous):

someone already helped me on this but idk where she got the numbers from

OpenStudy (anonymous):

this is the equation she solved find the perimeter of the shape

OpenStudy (anonymous):

@Skielerlucas04 can you help or do you know anyone who can

OpenStudy (anonymous):

@blackeistien238821 help please

OpenStudy (anonymous):

idk wut grade work dis

OpenStudy (anonymous):

geometry 10th grade

OpenStudy (anonymous):

i only 7th grade sorry

OpenStudy (anonymous):

wish i could help

OpenStudy (anonymous):

oh it thanks but do you know any that can help me

OpenStudy (anonymous):

jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

i think she got the numbers from the points but used slope intercep form to find the length of the sides

OpenStudy (anonymous):

so how did she get the numbers from slope intercept

jimthompson5910 (jim_thompson5910):

Focus on the right triangle I have drawn in red. This triangle has legs of 2 and 3.

jimthompson5910 (jim_thompson5910):

a = 2 b = 3 c = unknown a^2 + b^2 = c^2 2^2 + 3^2 = c^2 solve for c to get c = ???

OpenStudy (anonymous):

15

OpenStudy (anonymous):

sorry 12

jimthompson5910 (jim_thompson5910):

nope on both

OpenStudy (anonymous):

13

jimthompson5910 (jim_thompson5910):

2^2 + 3^2 = c^2 4 + 9 = c^2 c^2 = 13 c = ??

OpenStudy (anonymous):

what do i do with 13

OpenStudy (anonymous):

13 times 13 is 169

jimthompson5910 (jim_thompson5910):

the opposite actually, you take the square root of 13

jimthompson5910 (jim_thompson5910):

\[\Large c = \sqrt{13}\]

OpenStudy (anonymous):

that 3.605551275

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

The length of AB is \(\Large \sqrt{13}\) units long. The same reasoning applies to CD.

OpenStudy (anonymous):

or 4 lol

OpenStudy (anonymous):

i got 4 for CD

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

CD is the same length as AB

OpenStudy (anonymous):

but how did she get 19

jimthompson5910 (jim_thompson5910):

both are exactly sqrt(13) units long they are approximately 3.60555127546399 units long

jimthompson5910 (jim_thompson5910):

I don't see any 19

OpenStudy (anonymous):

like what IIIIIIIIIIIIIIIIII said

OpenStudy (anonymous):

12+2sqrt13 equals 19

jimthompson5910 (jim_thompson5910):

I'd have to see the full problem + instructions

OpenStudy (anonymous):

19 was the final answer

jimthompson5910 (jim_thompson5910):

Oh nvm, you're looking for the perimeter

jimthompson5910 (jim_thompson5910):

two sides are 6 6+6 = 12

jimthompson5910 (jim_thompson5910):

the other two sides are \[\Large \sqrt{13}\]

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{13}+\sqrt{13} = 2\sqrt{13}\]

OpenStudy (anonymous):

how are they 6

jimthompson5910 (jim_thompson5910):

Perimeter = \(\Large 12+2\sqrt{13} \approx 19.2111025509\) which rounds to 19

jimthompson5910 (jim_thompson5910):

BC = AD = 6

jimthompson5910 (jim_thompson5910):

you just count the spaces for the horizontal or vertical segments

OpenStudy (anonymous):

oh ok

jimthompson5910 (jim_thompson5910):



OpenStudy (anonymous):

were is 13 from

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

it's not just 13. It's the square root of it

OpenStudy (anonymous):

why cant it be sqrt 12 or another number why is it 13

jimthompson5910 (jim_thompson5910):

because that's what you get when you solve for c in 2^2 + 3^2 = c^2

jimthompson5910 (jim_thompson5910):

it's \(\sqrt{13}\) and NOT 13

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

@hershey21

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

idk lol